# Bridging and Modeling Correlations in Pairwise Data for Direct Preference Optimization

### v1

@@ Abstract
Direct preference optimization (DPO), a widely adopted offline preference optimization algorithm, aims to align large language models (LLMs) with human-desired behaviors using pairwise preference data. However, the winning response and the losing response within pairwise data are generated isolatedly, leading to weak correlations between them as well as suboptimal alignment performance. To address this issue, we propose an effective framework named BMC, for bridging and modeling correlations in pairwise data. Firstly, we increase the consistency and informativeness of the pairwise preference signals by targeted modifications, synthesizing a pseudo winning response through improving the losing response based on the winning response. Secondly, we identify that DPO alone is insufficient to model these correlations and capture nuanced variations. Therefore, we propose learning token-level correlations by dynamically leveraging the policy model’s confidence during training. Comprehensive experiments on QA, math, and instruction-following tasks demonstrate the effectiveness of our approach, significantly surpassing competitive baselines, including DPO. Additionally, our in-depth quantitative analysis reveals the reasons behind our method’s superior performance over DPO and showcases its versatility to other DPO variants. We release our repository at https://github.com/YJiangcm/BMC.
@@ 1 Introduction
Preference optimization is essential for aligning large language models (LLMs) with human values and intentions, e.g., helpfulness, honesty, and harmlessness (Askell et al., 2021; Ouyang et al., 2022). Reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Bai et al., 2022a; Ouyang et al., 2022) is a widely adopted technique for achieving this alignment effectively. Despite the impressive results of RLHF, it faces significant optimization challenges due to its multi-stage process, which includes training a reward model and subsequently optimizing a policy model to maximize this reward (Casper et al., 2023).
An effective and straightforward alternative to RLHF is direct preference optimization (DPO) (Rafailov et al., 2024b). Unlike the traditional RLHF approach, DPO bypasses training a reward model and avoids using any reinforcement learning algorithms. Instead, it reparameterizes the reward function to directly learn a policy model from offline pairwise preference data, employing the Bradley-Terry ranking objective (Bradley & Terry, 1952). However, the generation of winning and losing responses within preference data often occurs in an isolated manner, whether through human annotation (Bai et al., 2022a) or automated techniques such as RLAIF (Bai et al., 2022b) and reject sampling (Liu et al., 2023; Pace et al., 2024). This isolation implies that winning and losing responses are produced without mutual visibility, resulting in a lack of strong correlation or relevance between them. Consequently, this detachment hinders the capacity to discern subtle yet critical differences between responses, potentially leading to suboptimal optimization and model alignment.
@@
In this work, we introduce an innovative framework, termed BMC, to bridge and model correlations in pairwise data for direct preference optimization. During the Bridging Phase, we enhance correlations by increasing the consistency and informativeness of pairwise preference signals. By using the preferred response as a reference, we synthesize a pseudo preferred response through targeted modifications of the dispreferred response. This pseudo preferred response offers two key advantages: (1) it maintains high correlations with the dispreferred response while encapsulating all human-desired values in the preferred response; and (2) it serves as an interpolation between the winning and losing responses, easing the learning process. The nuanced differences between the pseudo-preferred and dispreferred responses are indeed what we expect the model to learn in the subsequent Modeling Phase. Nonetheless, we identify that DPO alone is insufficient to model these correlations and capture nuanced variations. From the perspective of the token-level Markov Decision Process (MDP) (Rafailov et al., 2024a), DPO aggregates rewards uniformly across all tokens, assuming equal contribution to sequence quality and neglecting token-specific importance. To address this, we adjust the emphasis on rewards of different tokens between the pseudo-preferred and dispreferred responses. Unlike previous methods (Guo et al., 2023; Cao et al., 2024; Chan et al., 2024; Chen et al., 2024a) that assign predefined values for fine-grained guidance, our adjustment is dynamically guided by the policy model’s confidence, i.e., the probability assigned to generated tokens during training. This ensures the model focuses on learning challenging distinctions while reinforcing known patterns, resulting in a more nuanced and robust policy.
@@
We conduct extensive experiments across three downstream scenarios: question answering, mathematical reasoning, and instruction following, utilizing a total of 10 datasets. Our results demonstrate that our method consistently and significantly outperforms competitive offline optimization algorithms across various tasks. Notably, our approach surpasses vanilla DPO by up to 3.8 points on question-answering tasks and by up to 6.4 points on AlpacaEval 2. Furthermore, we use in-depth analyses to elucidate why our method outperforms DPO and show that our framework can be versatilely adapted to other DPO variants, confirming its potential for broad application.
@@ 2 Related Work
Preference optimization refers to aligning large language models with human preferences or specific desired outcomes. A well-established method for this is reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Bai et al., 2022a; Ouyang et al., 2022), which typically involves three stages: supervised fine-tuning, reward model training, and policy optimization. Although RLHF produces highly capable models, its training process is often complex and unstable (Santacroce et al., 2023), requiring the training of multiple LMs and sampling from the LM policy. To address these challenges, Direct Preference Optimization (DPO) (Rafailov et al., 2024b) introduces an alternative offline algorithm to optimize the regularized expected rewards without relying on RL. Following the introduction of DPO, several studies have sought to enhance it in various ways. For instance, IPO (Azar et al., 2024) aims to prevent DPO from overfitting to the preference dataset; ORPO (Hong et al., 2024) and SimPO (Meng et al., 2024) eliminate the need for a reference model; and R-DPO (Park et al., 2024) incorporates a regularization term to prevent exploitation based on length.
@@
Constructing high-quality pairwise preference data is essential for preference optimization. Given the high cost of manually curating these datasets at scale, researchers have explored automated methods for producing preference data. One notable approach, RLAIF (Bai et al., 2022b) employs LLMs to label side-by-side response pairs, eliminating the need for human labeling. Alternatively, positive and negative responses can be generated by utilizing models of varying quality (Kim et al., 2023) or through specific prompting techniques (Yang et al., 2023). Recently, sampling-based methods such as Statistical Rejection Sampling (Liu et al., 2023) and West-of-N (Pace et al., 2024) have been introduced, generating preference pairs by selecting candidates sampled from the optimal policy. Nonetheless, these methods independently generate preferred and undesired responses without accounting for the correlations between them.
@@
The majority of preference optimization strategies typically utilize trajectory-wise (sequence-level) rewards, while LM training and generation both occur at the token level (Yang et al., 2024). To bridge this gap, FIGA (Guo et al., 2023) and DRLC (Cao et al., 2024) exploit external LLMs to pinpoint positive and negative token segments within responses, assigning fixed reward values (e.g., +1 for positive, -1 for negative) as guidance. Meanwhile, ABC (Chan et al., 2024) and RLMEC (Chen et al., 2024a) extract fine-grained credits from the reward model. Despite their contributions, these methods rely on predefined values for fine-grained guidance, failing to account for the dynamic learning process of the policy model.
@@ 3 Methodology
In this section, we present the proposed approach BMC by bridging and modeling correlations in pairwise data for direct preference optimization. As depicted in Figure 1, our BMC framework is structured around two pivotal stages: (1) the Bridging Phase, where we enhance the correlations between pairwise data by increasing the consistency and informativeness of pairwise preference signals through targeted modifications (§3.1); and (2) the Modeling Phase, where we dynamically model the correlations during the optimization process by leveraging the confidence of the policy model (§3.2), alleviating the insufficient token-level credit assignment of DPO.
@@ 3.1 Bridging Phase
In offline preference optimization, it is commonly assumed that we have access to a static pairwise preference dataset , where and denote the preferred and dispreferred completion, give the input prompt . To align an LLM, a basic principle is to ensure that the distribution of the model should not experience significant shifts during the alignment process (Bai et al., 2022a). Despite that is human preferred, it is likely to span a very different semantic distribution as . Additionally, since and are generated isolatedly, the correlation between and can be weak during pairwise preference optimization. This weak correlation poses a challenge, as the preferred outputs may not provide sufficiently informative gradients for adjusting the model’s parameters in relation to . Consequently, the optimization process may struggle to effectively leverage the provided pairwise preferences to fine-tune the model, potentially resulting in suboptimal alignment performance. To address this challenge, we enhance the alignment efficacy by improving the consistency and informativeness of the pairwise preference signals. Specifically, we utilize an off-the-shelf LLM to make targeted modification to by referring to :
where is the generated pseudo preferred completion, is the instruction prompt that requires to be modified only on dispreferred tokens compared to . In this way, we can effectively reduce the distribution shifts as well as obtain demonstrations more correlated to . Finally, we obtain a new dataset for subsequent training.
@@ 3.2 Modeling Phase
The difference between and can be easily obtained through dynamic programming algorithms like the Levenshtein Distance (Yujian & Bo, 2007), which highlights the core human expected and unexpected behaviors by decoupling from the inherent linguistic style and overall semantic distribution. To enhance the model’s ability to discern underlying patterns and deviations in human behavior, stopwords like “the” and “a” are filtered out, allowing for a more focused analysis of the substantive elements of the text. Despite these improvements, our findings indicate that DPO alone is insufficient for capturing the nuanced variances, suggesting the need for additional methods to fully model these correlations.
Limitations of DPO.
DPO (Rafailov et al., 2024b) stands as one of the leading methods for offline preference optimization. Rather than learning an explicit reward model, DPO reparameterizes the reward function using a closed-form expression with the optimal policy:
where and signify the policy model and reference model, respectively. is the partition function. By substituting the reparameterization into the Bradley-Terry model (Bradley & Terry, 1952) , DPO directly optimizes for the policy best satisfying the preferences with a simple classification objective:
DPO was originally derived as a bandit problem in which the whole response of the model is treated as a single arm. Alternatively, it can be deduced in the token-level MDP (Rafailov et al., 2024a):
@@
From Eq. 4, it can be seen that DPO assigns rewards to each token generation by , and simply add up the rewards of all tokens as the accumulated reward of the trajectory. This uniform aggregation assumes that each token contributes equally to the overall quality of the sequence, without considering the varying importance of each token (timestep). Therefore, nuanced differences between and that significantly influence the overall meaning or quality of the response might not be adequately emphasized (refer to Figure 6), leading to suboptimal performance.
@@
To this end, we propose to emphasize the rewards of critical tokens, i.e., nuanced differences between and . The magnitude of the emphasis is determined dynamically by the policy model’s confidence, which refers to the probability assigned to the generated token during training.
For varied tokens in the preferred response , low policy confidence signifies areas of insufficient learning, indicating a need for greater emphasis to enhance the model’s ability to discern these nuances. Consequently, we modify the reward factor to be inversely proportional to the policy model’s confidence, as demonstrated in Eq. 8.
For varied tokens in the dispreferred response , we adjust the reward factor by leveraging the autoregressive property of the policy model. Through the training of the policy model, it tends to assign a low probability to given the input . By analyzing token-level probabilities, we observe that the probability of the initial token in undesired spans is low, whereas the probabilities of subsequent tokens are significantly higher, as illustrated in Figure 2. This phenomenon arises from the autoregressive nature of the model, where subsequent mistakes are conditioned on previous errors. Although the onset of the undesired span should warrant penalization, the subsequent tokens might be contextually appropriate if the policy model’s confidence is high, justifying reduced penalization. Therefore, we modify the reward factor to be inversely proportional to the policy model’s confidence, as described in Eq. 11.
In a nutshell, the magnitude of the emphasis on critical tokens is determined dynamically by the policy model’s confidence. This adaptive mechanism ensures that the model focuses on learning challenging distinctions while reinforcing already learned patterns, leading to a more nuanced and robust policy. The formalization of this approach is encapsulated in Eq. 5, where and dynamically adjust based on the policy’s confidence, ensuring a tailored emphasis on critical tokens to improve the overall model performance.
@@
The denotes the stop-gradient operator, the is an upper limit threshold that controls the emphasis on the rewards of the critical tokens, preventing overly aggressive updates. The and signify using the Levenshtein Distance algorithm to find the varied tokens in and , respectively.
Gradient analysis.
For a mechanistic understanding of our method, we analyze the gradient of the loss function . The gradient with respect to the parameters can be written as:
where . It can be seen that our method takes into account both the sequential level and the token level perspectives: (1) at the sequential level, we encourage the preferred completion while punishing the dispreferred one; and (2) at the token level, we further refine the rewards of critical desired and undesired tokens of and , respectively. This dual consideration ensures that both the overall sequence structure and the individual token choices are optimized for the desired outcome.

### v3

@@ Abstract
Direct preference optimization (DPO), a widely adopted offline preference optimization algorithm, aims to align large language models (LLMs) with human-desired behaviors using pairwise preference data. However, the generation of the winning response and the losing response within pairwise data are typically isolated, leading to weak correlations between them as well as suboptimal alignment performance. To address this issue, we propose an effective framework for Bridging and Modeling Correlations in pairwise data, named BMC. Firstly, we increase the consistency and informativeness of the pairwise preference signals through targeted modifications, synthesizing a pseudo-winning response by improving the losing response with the winning response as a reference. Secondly, we identify that DPO alone is insufficient to model these correlations and capture nuanced variations. Therefore, we propose learning token-level correlations by dynamically leveraging the policy model’s confidence during training. Comprehensive experiments on QA, math, and instruction-following tasks demonstrate the effectiveness of our approach, significantly surpassing competitive baselines, including DPO. Additionally, our in-depth quantitative analysis reveals the reasons behind our method’s superior performance over DPO and showcases its versatility to other DPO variants. We release our repository at https://github.com/YJiangcm/BMC.
@@ 1 Introduction
Direct preference optimization (DPO) (Rafailov et al., 2024b) has emerged as a prominent alternative to reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Bai et al., 2022a; Ouyang et al., 2022) for aligning large language models (LLMs) with human values. Unlike the traditional RLHF approach, DPO bypasses training a reward model and avoids using any reinforcement learning algorithms. Since the inception of DPO, numerous studies have sought to advance this method by refining its training objective (Wang et al., 2024). For instance, IPO (Azar et al., 2024) introduces an alternative pairwise preference loss to mitigate overfitting to the preference dataset, while R-DPO (Park et al., 2024) incorporates a regularization term to prevent the exploitation of latent length bias in the training data.
However, relatively little attention has been given to enhancing DPO through advancements in the quality of preference data used for training. In particular, the generation of winning and losing responses within preference data often occurs in an isolated manner, either through human annotation (Bai et al., 2022a) or automated techniques such as RLAIF (Bai et al., 2022b) and reject sampling (Liu et al., 2024a; Pace et al., 2024). This isolation implies that winning and losing responses are produced without mutual visibility, resulting in a lack of strong correlation or relevance between them. Consequently, the model may struggle to identify nuanced yet significant distinctions that differentiate superior responses from inferior ones (Fürnkranz & Hüllermeier, 2010; Wirth et al., 2017), which can ultimately compromise optimization and alignment effectiveness.
@@
In this work, we introduce an innovative framework, termed BMC, to Bridge and Model Correlations in pairwise data for direct preference optimization. During the Bridging Phase, we enhance correlations by increasing the consistency and informativeness of pairwise preference signals. By using the winning response as a reference, we synthesize a pseudo-winning response through targeted modifications of the losing response. This pseudo-winning response offers two key advantages: (1) it preserves essential characteristics of the losing response, minimizing noise in preference signals (consistency); (2) it encapsulates all human-desired values from the winning response, enabling the model to better discern features that lead to superior performance (informativeness). The nuanced differences between the pseudo-winning and losing responses are indeed what we expect the model to learn in the subsequent Modeling Phase. Nonetheless, we identify that DPO alone is insufficient to model these correlations and capture nuanced variations. From the perspective of the token-level Markov Decision Process (MDP) (Rafailov et al., 2024a), DPO aggregates rewards uniformly across all tokens, assuming equal contribution to sequence quality and neglecting token-specific importance. To address this, we adjust the emphasis on rewards of different tokens between pseudo-winning and losing responses. Unlike previous methods (Guo et al., 2024; Cao et al., 2024; Chan et al., 2024; Chen et al., 2024a) that assign predefined values for fine-grained guidance, our adjustment is dynamically guided by the policy model’s confidence, i.e., the probability assigned to generated tokens during training. This ensures the model focuses on learning challenging distinctions while reinforcing known patterns, resulting in a more nuanced and robust policy.
@@
We conduct extensive experiments across three downstream scenarios: question answering, mathematical reasoning, and instruction following, utilizing a total of 10 datasets. Our results demonstrate that our method consistently and significantly outperforms competitive offline optimization algorithms across various tasks. Furthermore, we use in-depth analyses to elucidate why our method outperforms DPO and show that our framework can be versatilely adapted to other DPO variants, confirming its potential for broad application.
@@ 2 Related Work
Preference optimization refers to aligning large language models with human preferences or specific desired outcomes. A well-established method for this is reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Bai et al., 2022a; Ouyang et al., 2022). Although RLHF produces highly capable models, its training process is often complex and unstable (Santacroce et al., 2023). To address these challenges, direct preference optimization (DPO) (Rafailov et al., 2024b) introduces an alternative offline algorithm to optimize the regularized expected rewards without relying on RL. Building on DPO, subsequent methods like IPO(Azar et al., 2024) address overfitting to preference data, while R-DPO (Park et al., 2024) introduces length-based regularization to mitigate exploitation.
@@
Constructing high-quality pairwise preference data is essential for preference optimization. Given the high cost of manually curating these datasets at scale, researchers have explored automated methods for producing preference data. One notable approach, RLAIF (Bai et al., 2022b) employs LLMs to label side-by-side response pairs, eliminating the need for human labeling. Alternatively, winning and losing responses can be generated by utilizing models of varying quality (Kim et al., 2023) or through specific prompting techniques (Yang et al., 2023). Recently, sampling-based methods such as Statistical Rejection Sampling (Liu et al., 2024a) and West-of-N (Pace et al., 2024) have been introduced, generating preference pairs by selecting candidates sampled from the optimal policy. Nonetheless, these methods isolatedly generate winning and losing responses without accounting for the correlations between them.
@@
The majority of preference optimization strategies typically utilize trajectory-wise (sequence-level) rewards, while LM training and generation both occur at the token level (Yang et al., 2024). To bridge this gap, FIGA (Guo et al., 2024) and DRLC (Cao et al., 2024) exploit external LLMs to pinpoint positive and negative token segments within responses, assigning fixed reward values (e.g., +1 for positive, -1 for negative) as guidance. Meanwhile, ABC (Chan et al., 2024) and RLMEC (Chen et al., 2024a) extract fine-grained credits from the reward model. Despite their contributions, these methods rely on predefined values for fine-grained guidance, failing to account for the dynamic learning process of the policy model.
@@ 3 Methodology
In this section, we present the proposed BMC approach, which bridges and models correlations in pairwise data for direct preference optimization. As depicted in Figure 1, our BMC framework is structured around two pivotal stages: (1) the Bridging Phase, where we enhance the correlations between pairwise data by increasing the consistency and informativeness of pairwise preference signals through targeted modifications (§3.1); and (2) the Modeling Phase, where we dynamically model the correlations during the optimization process by leveraging the confidence of the policy model (§3.2), alleviating the insufficient token-level credit assignment of DPO.
@@ 3.1 Bridging Phase
In offline preference optimization, it is commonly assumed that we have access to a static pairwise preference dataset , where and denote the winning and losing response, give the input prompt . However, since and are typically generated in isolation, the correlation between and can be inherently weak during pairwise preference optimization. In the context of DPO, the Bradley-Terry objective (Bradley & Terry, 1952) computes gradients based on the relative likelihoods of and . When the correlation between and is weak, the differences between these responses are often superficial (e.g., stylistic or irrelevant variations) rather than substantive distinctions that reflect human-preferred behaviors. Consequently, the optimization process may inadvertently focus on minor discrepancies rather than meaningful distinctions. This results in gradients that are less informative for guiding the model towards robust preference alignment. To address this challenge, we enhance the alignment efficacy by improving the consistency and informativeness of pairwise preference signals. As shown in the upper part of Figure 1, we utilize an off-the-shelf LLM to make targeted modification of by referring to :
where is the generated pseudo-winning response, is the instruction (see examples in Appendix A.2) that requires to be modified only on dispreferred tokens, using as a reference guidance. In this way, preserves essential characteristics of the losing response while encapsulating all human-desired values in the winning response . The token-level differences between and highlight the core human expected and unexpected behaviors by decoupling from the inherent linguistic style and overall semantic distribution. Thus, refines the original training data for more focused learning, shifting the optimization process to concentrate on the most critical differences in preference data. The benefits of the Bridging Phase are further analyzed in §5.2. Finally, we use the new dataset for subsequent training.
An alternative approach that attempts to enhance the correlation between the winning and losing responses is to degenerate to via targeted modification and utilize as the preference pair. Nevertheless, our ablation study in Table 3 reveals that LLMs encounter challenges with this inverse operation, leading to a notable decline in performance.
@@ 3.2 Modeling Phase
After the Bridging Phase, the token-level differences between and can be obtained through dynamic programming algorithms like Levenshtein Distance (Yujian & Bo, 2007). As depicted in the lower part of Figure 1, these nuanced variations guide LLMs to prioritize the reinforcement of optimal actions while discouraging suboptimal ones within a single response. However, our findings below indicate that DPO alone is insufficient for capturing the nuanced variations, highlighting the necessity for supplementary techniques to comprehensively model these correlations.
Alternative interpretation of DPO.
DPO (Rafailov et al., 2024b) introduced a novel framework for optimizing the equivalent KL-constrained reward function as in RLHF, without the need to learn an explicit reward model. Instead, the problem is cast as a maximum likelihood estimation for the policy model on the preference dataset , resulting in the following training objective:
where is the reference model, typically the supervised fine-tuned (SFT) model, and is a regularisation term corresponding to the strength of KL-regularization in RLHF.
As shown in Eq. (2), DPO was originally conceptualized as a bandit problem, where the whole response of the model is treated as a single arm to receive a reward. More recently, Rafailov et al. (2024a) extended the theoretical foundation of DPO, showing that it can also be derived in the context of token-level MDP. The corresponding training objective at the token level is:
@@
It can be inferred from Eq. (3) that DPO, redefined as a token-level MDP, assigns rewards to each token generation by , and simply add up the rewards of all tokens as the accumulated reward of the trajectory. This uniform aggregation assumes that each token contributes equally to the overall quality of the sequence, without considering the varying importance of each token (timestep). Therefore, nuanced differences between and that significantly influence the overall meaning or quality of the response might not be adequately emphasized (refer to Figure 6), leading to suboptimal performance. To this end, we propose to emphasize the rewards of critical tokens, i.e., nuanced differences between and . The magnitude of the emphasis is determined dynamically by the policy model’s confidence, which refers to the probability assigned to the generated token during training. Below, we detail our design choices for the pseudo-winning response and losing response, respectively.
For varied tokens in the pseudo-winning response , we adapt the reward factor based on the learning process of the policy model. Lower policy confidence indicates underdeveloped learning of the target behavior, signaling the need for additional focus to help the model better capture these nuances. Consequently, we adjust the reward factor to be inversely proportional to the policy model’s confidence, as formalized in Eq. (7).
Figure 2: We aggregate varied tokens in or into more coarser-grained spans. During the DPO training on , we compute the averaged of tokens in different positions of spans. -
For varied tokens in the losing response , we carefully adjust the reward factor by reinforcing already learned patterns of the policy model. Intuitively, tokens in with higher confidence from the policy model may reflect inaccurate preference learning and therefore warrant stronger penalization. However, our analysis reveals a distinct pattern of the policy model when processing compared to . Specifically, when grouping varied tokens in into coarser-grained spans, the model’s confidence is significantly influenced by the token’s position within these spans, as illustrated in Figure 2. We observe that the probabilities assigned to the initial token of incorrect spans in are typically low, whereas the probabilities for subsequent tokens within the same span are notably higher. Prior studies have identified token probability as a critical signal for detecting anomalous behaviors (Xiao & Wang, 2021; Fadeeva et al., 2024) and assessing generation quality (Yuan et al., 2021; Fu et al., 2024). Consistent with these findings, our results indicate that during training, the policy model can effectively recognize the onset of undesired spans by assigning low probabilities to initial tokens. Nonetheless, due to the autoregressive dependencies, subsequent tokens within these spans receive higher probabilities, reflecting the contextual coherence established by preceding tokens, even when the span as a whole is incorrect. Thus, while it is crucial to penalize initial tokens, applying equally strong penalties to subsequent tokens might be suboptimal, as they often maintain local coherence within the flawed span. Therefore, we adjust the reward factor to also be inversely proportional to the policy model’s confidence in Eq. (10).
In a nutshell, our approach dynamically modulates the emphasis placed on critical tokens based on the policy model’s confidence. This adaptive reward mechanism ensures that the model focuses on learning challenging distinctions while reinforcing already learned patterns, ultimately fostering a more nuanced and robust policy (see our analysis in §5.2). The formalization of our approach is encapsulated in Eq. (4), where and adjust dynamically based on the policy’s confidence, ensuring a tailored emphasis on critical tokens to improve the overall model performance.
@@
The denotes the stop-gradient operator, the is an upper limit threshold that controls the emphasis on the rewards of the critical tokens, preventing overly aggressive updates. The and signify using the Levenshtein Distance algorithm to find the varied tokens in and , respectively. In Appendix C, we provide a gradient analysis of DPO-BMC. Unlike DPO, our approach harmonizes both sequential and token-level perspectives, effectively optimizing the overall sequence structure alongside crucial token choices for the desired outcome.